Love Econ shoes factory has the total cost function of 𝑇𝐶(𝑠)=500+3𝑠+5𝑠2.

Please complete the following:

(a) Total Variable Costs:

(b) Total Fixed Costs:

(c) Average Variable Costs:

(d) Average Fixed Cost:

(e) Average Total Costs:

(f) Marginal Costs:

(g) At what amount of output is average total cost minimized?

(h) Draw total fixed cost curve when Love is producing 300,000 units of shoes.


1
Expert's answer
2021-11-07T19:49:47-0500

Total Cost Function:

"TC = 500 + 3s + 5 s\\^2"

(a) Total Variable Costs:

 Fixed Costs= 500

Total variable cost includes only the variable component of the total cost.

Total Variable Costs: "3s + 5 s\\^2"

(b) Total Fixed Costs:

 Total fixed cost includes only the fixed parts of the total cost that does not change with the level of output.

Total Fixed Costs = 500

(c) Average Variable Costs: "\frac{Total\ Variable\ Costs}{s}"

"=\frac{3s + 5 s\\^2}{s}"

= "=3 + 5s"

(d) Average Fixed Cost:

"=\frac{Total\ Fixed\ Costs}{s}\\
=\frac{500}{s}"

(e) Average Total Costs:

Average Total Costs = "=\frac{Total\ Costs}{s}\\
\\
\frac{500 + 3s + 5 s\\^2}{s}\\
\frac{500 }{s} + 3 + 5s"

(f) Marginal Costs:

Marginal Costs = "\frac{dTotalCost}{ds}\ = 3 + 10s"

(g) At what amount of output is average total cost minimized?

Differentiating the average total cost with respect to s and equating it to 0 -

"\frac{-500}{s\\^2} + 5 = 0\\
s = 10"

Hence, output level = 10 the average total cost is minimized.

(h) Draw total fixed cost curve when Love is producing 300,000 units of shoes.